OpenStudy (wafels134):

Hey can anyone help me with The values in the table below were taken from the graph of a 3rd-degree polynomial function.

OpenStudy (wafels134):

The values in the table below were taken from the graph of a 3rd-degree polynomial function. x −3 −2 −1 0 2 P(x) 0 −20 0 30 0 Which of the following represents the equation of the polynomial function? P(x) = (x − 1)(x − 3)(x + 2) P(x) = −5(x + 1)(x + 3)(x − 2) P(x) = (x + 1)(x + 3)(x − 2) P(x) = 5(x + 1)(x + 3)(x − 2)

OpenStudy (phi):

They want you to know that if a polynomial is zero at x= a then (x-a) is a factor of the polynomial for example if we had x^2 -3x + 2 = 0 and we factor it into (x-1)(x-2) = 0 it should be clear that if x=1 then (x-1) becomes (1-1) = 0 and 0 * (1-2) = 0 or putting x=1 into the original x^2 -3x + 2 = 0 1 - 3*1 + 2 = 0 1 -3 +2 = 0 -2+2 = 0 0 = 0 we get zero (as we expected)

OpenStudy (phi):

In your problem you have x −3 −2 −1 0 2 P(x) 0 −20 0 30 0 and if we look only at the places where P(x) is 0 x −3 −1 2 P(x) 0 0 0 we have 3 places: x= -3 , x= -1 and x= 2 that means the 3 factors are (x- -3), (x- -1) and (x- 2) they simplify (using minus a minus is plus) to (x+3)(x+1)(x-2) so the correct answer must have those 3 factors.

OpenStudy (phi):

that means we can cross off the first choice. 3 choices remain. to figure out which of the 3 is correct, I would use this info x=0, P(x) = 30 (when x is 0, we should get 30) so in each of the 3 choices, replace x with 0 and simplify if you don't get 30, cross off that choice.